answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of rolling a number greater than 3 on one roll of a die?

Sample space, roll of 1 die, is: 1, 2, 3, 4, 5, 6. The numbers greater than 3 are: 4, 5, 6; which is 1/2 of the sample space. So, probability of rolling a number greater than 3 on one roll of a die is 1/2 or 0.5.

If a 6 sided number cube is rolled twice how many possible outcomes are there?

6*6=36

1&1; 1&2; 1&3; 1&4; 1&5; 1&6

2&1; 2&2; 2&3; 2&4; 2&5; 2&6

3&1; 3&2; 3&3; 3&4; 3&5; 3&6

4&1; 4&2; 4&3; 4&4; 4&5; 4&6

5&1; 5&2; 5&3; 5&4; 5&5; 5&6

6&1; 6&2; 6&3; 6&4; 6&5; 6&6

For adding up roll one and roll two, the outcomes are

1+1= 2

1+2=2+1= 3

1+3=2+2=3+1= 4

1+4=2+3=3+2=4+1= 5

1+5=2+4=3+3=4+2=5+1= 6

1+6=2+5=3+4=4+3=5+2=6+1= 7

2+6=3+5=4+4=5+3=6+2= 8

3+6=4+5=5+4=6+3= 9

4+6=5+5=6+4= 10

5+6=6+5= 11

6+6= 12

Odds of adding up to those possible outcomes.

2 = 1:36

3 = 2:36 = 1:18

4 = 3:36 = 1:12

5 = 4:36 = 1:9

6 = 5:36

7 = 6:36 = 1:6 You will total seven more times than others.

8 = 5:36

9 = 4:36 = 1:9

10= 3:36 = 1:12

11= 2:36 = 1:18

12= 1:36

In rolling two fair dice what is the probability of obtaining a sum greater than 10 or a sum divisible by 6?

In rolling two fair dice, the probability of obtaining a sum greater than 10 is 3 in 36 or 1 in 12, and the probability of obtaining a sum divisible by 6 is 6 in 36 or 1 in 6. Since the two probabilities come from the same population of 36 possible outcomes, the combined probability is the inclusive or of the two result sets, 8 in 36 or 2 in 9.

The results that match the sum>10 target are 5-6, 6-5, and 6-6.

The results that match the divisibility by 6 target are 1-5, 2-4, 3-3, 4-2, 5-1, and 6-6.

The results that match the combined target are 1-5, 2-4, 3-3, 4-2, 5-1, 5-6, 6-5, and 6-6.

If you roll a dice 100 times how many times would you roll a two?

There are six possible outcomes. Assuming the probability of each outcome is the same (dice has no defects), then you are likely to roll the number two, 100/6=50/3=16.67 times.

Probability to get exactly 2 heads when flipping a coin 3 times?

The probability of throwing exactly 2 heads in three flips of a coin is 3 in 8, or 0.375.

There are 8 outcomes of flipping a coin 3 times, HHH, HHT, HTH, HTT, THH, THT, TTH, and TTT. Of those outcomes, 3 contain two heads, so the answer is 3 in 8.

What is the probability of two dice that have a 1 side with a one 2 sides with two on it and 3 sides with 3 on it. and not getting a 2 when you role both?

First you need to figure out the probability of it not landing on a 2 when you roll one die. Since there are six sides and two sides with 2 on it, the probability that one die will not land on 2 is 4/6.

Now you multiply 4/6 by 4/6 to find the correct odds as to both die not landing on 2.

4/6 * 4/6 = 16/36 = 4/9

The percentage is (4/9)*100% = (approx.) 44.4%

What is the exponential growth formula?

y = a * bxUsing "^" for powers. Any exponential formula can be written in one of the following two forms:

y = A e^Bx

y = C D^x

where "y" is the independent variable, "x" is the dependent variable, and "A", "B", "C", and "D" are arbitrary constants.

Find the probability that the month of February in a year 2008 may have 5 Tuesdays?

Considering the fact that a non-leap year having 28days in February will definitely have 4 Tuesdays , we have to find the probability that the 29th day of a leap year (2008) is also a Tuesday. This probability is 1/7 i.e 1 out the probable 7days in a week !

Well this said, this question is void in a sense knowing that we are in 2010 and that February 2008 had 5 Fridays and therefore the probability of 5 Tuesdays is ZERO as of 2008 !

How do you find area under a curve between two z values?

Let z1 be the smaller z value and z2 be the larger value. Let N(z) be the cumulative normal distribution evaluated at some value z. The random variable is Z.

The probability that Z has a value from minus infinity to z2 is equal to N(z2). You can show this by drawing a bell shape curve, and shading in everything to the left of z2 as equal to the area under curve.

Similarly, the probability that Z has a value from minus infinity to z1 is N(z1).

The area under the bell curve (standard normal cumulative distribution) is N(z1) - N(z2). I can show this with a little example:

z1= -1 z2 = 2 Area = N(2) - N(-1) = 0.9773 - 0.1587 = 0.8186. I used Excel with the normsdist(z) function. The mean is zero and standard deviation is one with this function.

What does a low mean absolute deviation mean?

It means that the observations are all close to their mean value.

Use the line of random numbers below to select the first six students to receive the treatment What is the name of the fifth student selected?

A psychologist wishes to conduct a study on the effects of music deprivation on high school students. A high school class consists of the 30 students numbered in the list below. The researcher establishes a treatment group of 15 students who will have their portable music players replaced by experimental players that present the sound of water running. The control group of 15 students will all get regular portable music players, stuffed full of their favorite songs.

Use the line of random numbers below to select the first six students to receive the treatment. What is the name of the fifth student selected?

59784 44312 15954 09233 00046 74318 02610 57396 16843 38454

Quincy

What is the probability of 26 people in a group having the same birthday?

Birthdays are not uniformly distributed over the year. But, if you do assume that they are, then:

ignoring leap years, it is approx 0.5982.

If you include leap years it is 0.5687.

Two dices are thrown simultaneously what is the probability that we get an even number on 1 of the dice?

If two dice (not dices!) are thrown then the probability that you get an even number on one die is 27/36 = 3/4

What does a negative z score mean?

It means that it is less than the mean.

What is the probability that a group of five random digits will contain at least 3 even digits?

Use the entire excerpt of your printable table of random numbers, shown below, to estimate the answer to this question: What is the probability that a group of five random digits will contain at least three even digits? (Zero is considered an even digit.)

46370 55170 53480 49126 89212 75292 67291 88241 37808 38154

Express your answer as a fraction using the slash ( / ) bar. Reduce fractions to their lowest terms.

9/10

Trending Questions
If you have 5 blue marbles 3 red marbles and 7 yellow marbles what is the probability of not getting a yellow if you draw one of them out of a sack? What exact percent falls the normal curve falls below at Z score of -2.00? How do you get 1000 stickers on first in math? How many cards are not a club in a 52 deck? For statistics class I Have to find the probability that exactly 41 murders committed? Is homogeneous uniform distribution? What is the distribution of sample means constructed by sampling 5 items from a population of 15? What is true of a random sample? Why did the two florists get married probability worksheet? A box contains identical cards on each card a letter of the English alphabet is printed.A card is selected at random from the box and not replaced and a second card is then selected. Find the probability that two cards selected(I) have vowels printed? Give the reasons why sample are prefered to entire population? What are the probabilities of getting a 7? What are the odds of rolling 6 consecutive doubles with 2 dice? What is the probability of rolling a whole number on one roll of a number cube? What is the probability of the next four births at a hospital being all girls? What is a probability based upon an event that has already occurred? If a coin is tossed 3 times what is the probability of getting 2 and only 2 heads? How does one find the probability of A and B compliment? Probability of a tsunami on the east coast of the US? One letter is randomly selected from the word math and a second letter is randomly selected from the word jokes what is the probability that both letters are vowels?